OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that is parallel to y=-3x+2 and passes through (2, -6). help please (: <3

OpenStudy (aroub):

y=mx+b -6=-3(2)+b -6=-6+b b=0 y=-3x

OpenStudy (anonymous):

YOUR AMAZING (: thank you sooooo much!!

OpenStudy (aroub):

hahaha anytime :)

OpenStudy (anonymous):

(: